The high-speed winds around a tornado can drive projectiles into trees, building walls, and even metal traffic signs. In a laboratory simulation, a standard wood toothpick was shot by pneumatic gun into an oak branch. The toothpick mass was 0.14 g, its speed before entering the branch was 228 m/s, and its penetration depth was 17 mm. If its speed was decreased at a uniform rate, what was the magnitude of the force of the branch on the toothpick?
Answer #1

mass = m = 0.14 x 10^-3 kg

initial velocity = V0 = 228 m/sec

final velocity = V = 0 m/sec

distance = S = 17 mm = 0.017 m

V^2 - V0^2 = 2 a S

0 - ( 228)^2 = 2 * a * 0.017

acceleration = a = 15.3 x 10^5 m/sec^2

force = F = m a = 0.14 x 10^-3 * 15.3 x 10^5 = 214.2 N

force = F = 214 N